OpenStudy (anonymous):

Given the function g(x) = 8x - 2, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. Answer (A)The value of g(-2) is larger than the value of g(4). (B)The value of g(-2) is smaller than the value of g(4). (C)The values of g(-2) and g(4) cannot be compared. (D)The value of g(-2) is the same as the value of g(4)

OpenStudy (anonymous):

g(x)=8x-2, g(-2)=8(-2)-2=-16-2=-18 g(4)=8(4)-2=32-2=30 g(4)>g(-2)

OpenStudy (anonymous):

can you please explain to me how you just simplified that equation, i never was good with algebraic equations.